OpenStudy (anonymous):

Use the order of operations to simplify this expression: [3 + 4(5 · 22) – 8 / 2].

OpenStudy (anonymous):

I get 439

OpenStudy (anonymous):

can you please show me how you got that answer

OpenStudy (anonymous):

You have PEMDAS which is Parenthesis Exponents Multiplication Division Addition Subtraction which is the order of operations so for parenthesis you have 5*22 which equals 110 then you go across and look for Multiplication since you do not have any exponents and you get 4*110 which equals 440 Next you look for division and get -8/2 which equals -4 So now you have the equation 3+440-4 440+3=443 and 443-4=439

OpenStudy (anonymous):

439 is correct. First you multiply 5 and 22. This comes to 110. Then you multiply the 4 and 110. This gives you 440. The next step is division. 8/2 = 4. Then addition. 440+3=443. subtract 4, you get 439.

OpenStudy (anonymous):

the correct equation is [3 + 4(5 · 2^2) – 8 / 2]?

OpenStudy (anonymous):

oh wait, its not 22 its 2 ^2

OpenStudy (anonymous):

oh then the answer is 79 3+4(5*4)-8/2 3+4(20)-8/2 3+80-8/2 3+80-4 83-4 79

OpenStudy (anonymous):

Thank you so much for your guys help :)